Reports 1-1 of 1 Clear search Modify search
AdV-PSL (Laser Bench setup)
turconi, soulard, chaibi, chiummo - 17:18 Friday 16 October 2020 (49676) Print this report
Beam dump on the slave laser

Yesterday afternoon we blocked the beam of the slave laser. Two beam dumps were put on the beam path (pointed by the arrows in the picture). The big one blocks the SL beam, the small one blocks the ML beam. Both dumpers are not clamped on the bench. --> Thus the SL is not locked, the signals PSL_ML_DC , PSL_ML_AC and PSL_SL_REFL_PD1_DC are not at their nominal value. 

This activity is part of the investigations on fast unlocks. We are now sure that no feedback light from the slave goes into the master. We recall that we are now monitoring the feedback light on the signal PSL_ML_FREQ_Q_MONIT_FS . Since yesterday 13:45 UTC we added an amplifier x10 on this signal.

The IMC and RFC were relocked yesterday afternoon at about 14:00 UTC and no unlocks were observed last night.

Images attached to this report
Comments to this report:
mwas - 15:01 Sunday 18 October 2020 (49683) Print this report

Note that there are still times when the noise on the EOM is increased from +/-0.1V to +/-0.5V see figure 1 and 2. 

Thes times correspond to a broad band (up to 200kHz or 10kHz) increase in noise on the ML PZT, the frequency monitor and the EOM correction.

Images attached to this comment
Search Help
×

Warning

×